OpenStudy (miranda090701):

Use number properties to simplify the following expression. -5 + (5 + 3) In the box below, show each step in simplifying the expression and explain which property you used in each step.

OpenStudy (solomonzelman):

i don't even know how to start :::::

OpenStudy (solomonzelman):

Parenthesis first

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \displaystyle \rm P }\)arenthesis \(\Huge\color{green}{ \displaystyle \rm E }\)xponents \(\Huge\color{orangered}{ \displaystyle \rm M }\)ultiply \(\Huge\color{orangered}{ \displaystyle \rm D }\)ivide \(\Huge\color{darkgoldenrod}{ \displaystyle \rm A }\)dd \(\Huge\color{darkgoldenrod}{ \displaystyle \rm S }\)ubtract

OpenStudy (solomonzelman):

this is the abbreviation for the order of operations....

OpenStudy (solomonzelman):

So first you do the "5+3" in the parenthesis, what is 5+3?

OpenStudy (miranda090701):

It's talking about the properties not operation

OpenStudy (solomonzelman):

well, you say say \(\large\color{black}{-5+(\color{red}{5+3})~~~\Rightarrow~~~\underline{-5+5}+3 }\) ( \(\large\color{black}{\Uparrow }\) you do this first)

OpenStudy (solomonzelman):

same as -b+(b-a) = -b+b +a=0+a=a

OpenStudy (miranda090701):

okay

OpenStudy (miranda090701):

thanks!!!!

OpenStudy (solomonzelman):

yw